OpenStudy (candy13106):

20n-2=10n+8 solve the equation

OpenStudy (anonymous):

20n−2=10n+8(Simplify both sides of the equation) 20n−2+−10n=10n+8+−10n(Add -10n to both sides) 10n−2=8 10n−2+2=8+2(Add 2 to both sides) 10n=10 (10n) 10 = 10 10 (Divide both sides by 10) n=1

OpenStudy (anonymous):

Please click as best response if this helped. :)

OpenStudy (candy13106):

huh

OpenStudy (arabpride):

20n - 2 = 10n + 8 -10n -10n 10n - 2 = + 8 +2 +2 10n = 10 /10 10 n = 1

OpenStudy (anonymous):

n = 1

OpenStudy (anonymous):

Told ya.. lol

OpenStudy (anonymous):

Answer: n = 1

OpenStudy (arabpride):

Techinically mine was the best because i explained so :P lol

OpenStudy (anonymous):

So did I. I broke it down

OpenStudy (candy13106):

thanks to both

OpenStudy (anonymous):

No problem.

OpenStudy (candy13106):

yep

OpenStudy (anonymous):

As long as you learn. :)

OpenStudy (arabpride):

WElcome *^.^* Any Time

OpenStudy (arabpride):

ooooooh i have an idea: @Candy13106 gives me a medal I give it to @Nikkeyiboo and @Nikkeyiboo givez it to @Candy13106

OpenStudy (arabpride):

So that way we all get medal *^.^*